Zelman has more than eighteen years his hijo.hace a year, she
had three times the age of the boy. that age now?

Answer by Alan3354(69443)

Zelman has more than eighteen years his hijo.hace a year, she
had three times the age of the boy. that age now?
-------------------
If I understand this, Zelman is 18 years older than her son. In one year, she will be 3 times his age.
----------------------
Z = Zelman's age
S = son's age
--------------
Z = S+18
Z+1 = 3*(S+1)
--------------
Z+1 = 3S+3
Sub S+18 for Z
S+18 + 1 = 3S+3
S+19 = 3S+3
-2S = -16
S = 8, the age of el hijo.
Z = S+18 = 26, Zelman's age.